# Realistic assessment of a single gravitational wave source localization taking into account precise pulsar distances with pulsar timing arrays

### v1

@@ I Introduction
Similar to the stochastic GW background from the SMBHBs, the detection of continuous gravitational waves (CGWs) from individual SMBHBs will also be expected [14, 15]. Among the information obtained by detecting the CGWs, information about the sky location of the GW source is important to identify its host galaxy. Numerous papers have studied the CGWs with mention of the sky location of the GW source [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50]. For example, in the CGW analysis of the IPTA Second Data Release [46], there is a case where the sky location of the GW source is relatively well constrained, although the uncertainty remains on the order of 100 to 1000 square degrees. It has also been suggested that the angular separation between the pulsars and the sky location of the CGW signal model is an important factor in interpreting these localization results.
In our previous study [47], the effect of precise pulsar distance measurements on the determination of sky location of a single GW source is investigated, and it was found that a few pulsars with accurately measured distances can improve the sky localization by several orders of magnitude. In that analysis, we assumed that the pulsars were uniformly distributed on the celestial sphere. However, in reality, pulsars are predominantly concentrated along the Galactic plane. Based on the IPTA CGW analysis results, the localization accuracy of the CGW source is likely to depend on the relative spatial configuration between the pulsars and the source. Therefore, in this study, we consider the actual spatial distribution of the pulsars used in current pulsar timing arrays.
@@ II Continuous GW Signal
where ${\hat{x}},{\hat{y}}$ and ${\hat{z}}$ are the basis vectors in the SSB Cartesian coordinate system and $\hat{\Omega}=\hat{m}\times\hat{n}$. The sky location of the GW source is specified by the polar angle $\theta$ and the azimuthal angle $\phi$. The antenna pattern function of the PTA, which encodes the angular sensitivity of the GW detector, is given by [55]:
$F^{A}_{p}({\hat{\Omega}})={\frac{1}{2}}{\frac{{\hat{p}_{p}}^{\mu}{\hat{p}_{p}}^{\nu}}{1+{\hat{\Omega}}\cdot{\hat{p}_{p}}}}e_{\mu\nu}^{A}({\hat{\Omega}}),$ (7)
@@ III Simulated Data
[…]
No evolution - - 1 $10^{8.2}$ $M_{\odot}$ $10^{-8}$ ${\rm Hz}$ $10^{0.67}$ ${\rm Mpc}$ 1 1 1 kpc $44$ ${\rm ns}$ 951650 yr
Slow evolution - - 1 $10^{9}$ $M_{\odot}$ $10^{-8}$ ${\rm Hz}$ $10^{2}$ ${\rm Mpc}$ 1 1 1 kpc $44$ ${\rm ns}$ 44172 yr
Fast evolution - - 1 $10^{9.8}$ $M_{\odot}$ $10^{-8}$ ${\rm Hz}$ $10^{3.33}$ ${\rm Mpc}$ 1 1 1 kpc $44$ ${\rm ns}$ 2050 yr
@@ IV.1 Parameter Estimation
where $s_{p}(t_{a},\hat{\Omega})$ of Eq. 9 is observed at a specific TOA, and $a=1,\dots,N_{\rm TOA}$. We also define that all of the noise for all of the pulsars and TOAs are characterized by the same Gaussian white noise.
@@ IV.2 Closed Form Posterior
When the noise is sufficiently small compared to the signal, the posterior distribution is expected to be sharply peaked around the true value of the parameter. Follwing articles [59, 60], we expand the CGW signal model $\bm{s}(\vec{\theta})$ to first order around the true parameter value $\vec{\theta}_{0}$ as
@@
where index $i$ denotes the component of the vector for the parameter and the Einstein summation convention is used for Latin indices throughout this section. Follwing articles [59, 60], we refer to this approximation as the high signal-to-noise ratio / linearized-signal approximation (high-SNR/LSA approximation). To compute the above function, the true values of the CGW signal model must be known. However, even if the true values are known, the shift of the peak of the posterior distribution from the true values due to noise remains unknown.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ IV.4 Constant Probability Density Contour
Using the closed-form posterior distribution, which is the multivariate normal distribution obtained in Section IV.2, the constant probability density contour can be calculated [62]. The definition of the ellipsoidal contour of the multivariate normal distribution is given in Appendix A.
@@ V Result
First, we present the results of estimating the localization accuracy of a GW source in the absence of precise pulsar distance information. In this case, we assume an uncertainty of 100 pc for all pulsar distances, which corresponds to the fiducial case described in Table 3. Since this uncertainty is much larger than the GW wavelength ($\sim$ 1 pc) it provides little information for determining the phase of the pulsar term. Fig. 2 shows the localization accuracy of GW sources as a function of the GW sky location under white noise levels of 10, 30, and 100 ns. In all cases, the accuracy is higher (the area is smaller) in regions where pulsars are densely distributed. It can be seen that even in the region with the best localization accuracy, $10^{-3}$ $\rm deg^{2}$ cannot be achieved. More details regarding the dependence of localization area on white noise level are provided in Appendix B.
@@
Figure 5: Left panel: The source numbers corresponding to the contours in Fig. 4. Right panel: Areas $\Omega$ of the two-dimensional 68% contours for GW sources located between J0437-4715 and J0030+0451. Source 14 is the same as Source 5. The green lines correspond to the fiducial case, while the blue lines correspond to Case 3. The solid, dashed, and dash-dotted lines correspond to the Slow, No, and Fast evolution cases in Table 1, respectively.
@@ VI Conclusions and Discussions
There are two major limitations in the current analyses that evaluate parameter estimation capability using the closed-form posterior distribution. The first is the assumption that the noise is sufficiently small for the CGW signal to be described by a linearized model, referred to as the high-SNR/LSA approximation, resulting in the posterior distribution with a single peak. However, current CGW analyses exhibit multimodal structure in sky location [46], which cannot be captured by our approach, while this may not be the case in the SKA era. The second limitation is that, although we performed Gaussian integration to obtain the closed-form expression, the parameters of stochastic signals cannot be treated as free parameters. Investigating the behavior of simultaneous parameter estimation of a CGW and stochastic signals remains a subject for future work. In particular, understanding the simultaneous parameter estimation of common-spectrum spatially uncorrelated red noise and a CGW, as demonstrated in [12], is likely to become increasingly important. Recent studies also emphasized the importance of resolving a CGW and the stochastic GW background simultaneously [48, 50].
@@ Acknowledgements.
RK is supported by JSPS KAKENHI Grant Number 24K17051. KT is partially supported by JSPS KAKENHI Grant Numbers 20H00180, 21H01130, 21H04467, and 24H01813, Bilateral Joint Research Projects of JSPS, and the ISM Cooperative Research Program (2023-ISMCRP-2046).
Appendix A Contour of the multivariate normal distribution
@@
Substituting Eq. 52 into Eq. 51 shows that the probability density on the surface of the ellipsoid is independent of $\vec{x}$ and is constant. This surface is called a contour. According to Eq. 52, it is represented as an ellipsoid centered at $\vec{\mu}$, with axes given by ±$\pm\sqrt{\chi^{2}_{k}(\alpha)\lambda_{i}}\vec{e}_{i}$, where ($\lambda_{i}$, $\vec{e}_{i}$) is an eigenvalue-eigenvector pair of $\Sigma$, which satisfies $\Sigma\vec{e}_{i}=\lambda_{i}\vec{e}_{i}$ for $i=1,2,\cdots,k$.
Appendix B Dependence of area on white noise level
Fig. 6 shows how the estimated area depends on the white noise level. It can be seen that as the white noise level decreases, the result approaches that of the uniform prior case. Conversely, as the noise level increases, it approaches that of the fixed parameter case. The intermediate region between the uniform prior case and the fixed-parameter case is where the Gaussian prior exhibits nontrivial behavior. In the fiducial case, the difference between the Gaussian prior and the uniform prior becomes apparent when the white noise exceeds 10 ns. It should be noted that when the white noise exceeds the amplitude of the Earth term, the high-SNR/LSA approximation underlying the closed-form posterior distribution becomes invalid, leading the resulting posterior distribution to deviate from the true posterior distribution.
$\propto\frac{1}{\sigma_{n}},$ (54)
where $\vec{\theta}_{\rm ML}$ is the maximum likelihood estimator. This relationship is consistent with the result of a frequentist study [18]. It is known that, under the high-SNR/LSA approximation, the covariance matrix of the posterior distribution with a uniform prior, as shown in Eq. 39, coincides with that of the maximum likelihood estimator [60]. Since the area is characterized by the covariance matrix, this consistency is an expected result. It is also known that the mean of the maximum likelihood estimator equals the true value. Therefore, in Fig. 6, we approximated the SNR by substituting the true value in place of the estimator itself. Although the difference between the maximum likelihood estimator and its mean is expected to increase with larger white noise, this effect is beyond the scope of the present analysis.

### v2

@@ I Introduction
Similar to the stochastic GW background from the SMBHBs, the detection of continuous gravitational waves (CGWs) from individual SMBHBs will also be expected [50, 37]. Among the information obtained by detecting the CGWs, information about the sky location of the GW source is important to identify its host galaxy. Numerous papers have studied the CGWs with mention of the sky location of the GW source [17, 51, 52, 40, 8, 20, 21, 46, 59, 64, 70, 65, 71, 73, 7, 33, 41, 58, 72, 67, 66, 9, 30, 44, 12, 60, 54, 63, 11, 16, 23, 36, 5, 13, 24, 38, 39, 47, 28, 29]. For example, in the CGW analysis of the IPTA Second Data Release [23], there is a case where the sky location of the GW source is relatively well constrained, although the uncertainty remains on the order of 100 to 1000 square degrees. It has also been suggested that the angular separation between the pulsars and the sky location of the CGW signal model is an important factor in interpreting these localization results.
In our previous study [36], the effect of precise pulsar distance measurements on the determination of sky location of a single GW source is investigated, and it was found that a few pulsars with accurately measured distances can improve the sky localization by several orders of magnitude. In that analysis, we assumed that the pulsars were uniformly distributed on the celestial sphere. However, in reality, pulsars are predominantly concentrated along the Galactic plane. Based on the IPTA CGW analysis results, the localization accuracy of the CGW source is likely to depend on the relative spatial configuration between the pulsars and the source. Recent simulation studies have pointed out that, to improve the localization accuracy, a small angular separation between the CGW source and nearby pulsars plays a significant role [47]. Therefore, in this study, we consider the actual spatial distribution of the pulsars used in current pulsar timing arrays.
@@ II Continuous GW Signal
where $^{\hat{x}},{\hat{y}}$ and $^{\hat{z}}$ are the basis vectors in the SSB Cartesian coordinate system and $\hat{\Omega}=\hat{m}\times\hat{n}$. The sky location of the GW source is specified by the polar angle $\theta$ and the azimuthal angle $\phi$. The antenna pattern function of the PTA, which encodes the angular sensitivity of the GW detector, is given by [14]:
[…]
@@ III Simulated Data
$\theta$ [^1] We consider various values for GW sky locations ($\theta$ and $\phi$). $\phi$11footnotemark: 1 $\iota$ $\mathcal{M}$ $f_{0}$ $d_{L}$ $\Phi_{0}$ $\psi$ $L_{p}$ $A_{e}$ $t_{\rm coal}$
No evolution - - 1 $10^{8.2}$ $M_{\odot}$ $10^{-8}$ $\rm Hz$} $10^{0.67}$ $\rm Mpc$} 1 1 1 kpc 4444 $\rm ns$} 951650 yr
Slow evolution - - 1 $10^{9}$ $M_{\odot}$ $10^{-8}$ $\rm Hz$} $10^{2}$ $\rm Mpc$} 1 1 1 kpc 4444 $\rm ns$} 44172 yr
Fast evolution - - 1 $10^{9.8}$ $M_{\odot}$ $10^{-8}$ $\rm Hz$} $10^{3.33}$ $\rm Mpc$} 1 1 1 kpc 4444 $\rm ns$} 2050 yr
Finally, we discuss the amplitudes of CGW signals that are currently detectable. For the slow-evolution case, we adopt $\mathcal{M}=10^{9}\,M_{\odot}$, $f_{0}=10^{-8}\,{\rm Hz}$, and $d_{L}=10^{2}\,{\rm Mpc}$. For the same mass and frequency, current PTA limits place a lower bound on the distance of approximately 10 Mpc [5]. Consequently, the slow-evolution case corresponds to a signal whose amplitude is about one-tenth of the current observational limit, since the amplitude of a CGW signal scales inversely with distance, as given by Eqs. 13 and 14. Thus, the amplitude is below the detection threshold of current PTAs. Nevertheless, it is sufficiently large to be considered realistic for simulations of future high-precision PTA datasets.
@@ IV.1 Parameter Estimation
[…]
@@ IV.2 Closed Form Posterior
In the high signal-to-noise ratio (SNR) regime, the posterior distribution is expected to be sharply peaked around the true value of the parameter. Follwing articles [62, 49], we expand the CGW signal model $\bm{s}(\vec{\theta})$ to first order around the true parameter value $\vec{\theta}_{0}$ as
@@
where index $i$ denotes the component of the vector for the parameter and the Einstein summation convention is used for Latin indices throughout this section. The previous article has shown that the posterior distribution obtained using a linearized waveform corresponds to the leading-order term in an expansion of the posterior distribution in powers of the inverse of the optimal SNR of the true signal [62]. In this context, the optimal SNR of the true signal represents the total SNR of the dataset, as defined in Eq. 51 in Appendix, but it is evaluated using the true parameter values of the CGW signal model rather than the maximum-likelihood estimates. Follwing articles [62, 49], we refer to this limit as the high signal-to-noise ratio / linearized-signal approximation (high-SNR/LSA) limit. To compute the above function, the true values of the CGW signal model must be known. However, even if the true values are known, the shift of the peak of the posterior distribution from the true values due to noise remains unknown.
@@
$\int d\vec{\theta}\exp\left[-\frac{1}{2}\left(\theta_{i}A_{ij}\theta_{j}-2b_{i}\theta_{i}\right)\right]=\sqrt{\det(2\pi A^{-1})}\exp\left[\frac{1}{2}b_{i}\left(A^{-1}\right)_{ij}b_{j}\right],$ (35)
@@
$\langle f(\vec{\theta})\rangle\equiv\int d\vec{\theta}\,f(\vec{\theta})p(\vec{\theta}|\bm{\delta t},M).$ (40)
@@
$p(\vec{\theta}|M)$ $=\frac{1}{\sqrt{\det(2\pi D)}}\exp\left[-\frac{1}{2}\left(\theta_{i}-\theta_{0i}\right)\left(D^{-1}\right)_{ij}\left(\theta_{j}-\theta_{0j}\right)\right]\quad\text{(Gaussian prior)}.$ (41)
@@
$\Sigma_{ij}$ $=\left((H+D^{-1})^{-1}\right)_{ij}\quad\text{(Gaussian prior)}.$ (44)
@@
$\langle f({\bm{\delta t}})\rangle_{n}$ $\equiv\int d{\bm{\delta t}}\,f({\bm{\delta t}})p({\bm{\delta t}}|\vec{\theta}_{0},M),$ (45)
@@ IV.4 Constant Probability Density Contour
Using the closed-form posterior distribution, which is the multivariate normal distribution obtained in Section IV.2, the constant probability density contour can be calculated [35]. The definition of the ellipsoidal contour of the multivariate normal distribution is given in Appendix B.
@@ V Result
First, we present the results of estimating the localization accuracy of a GW source in the absence of precise pulsar distance information. In this case, we assume an uncertainty of 100 pc for all pulsar distances, which corresponds to the fiducial case described in Table 3. Since this uncertainty is much larger than the GW wavelength ($\sim$ 1 pc) it provides little information for determining the phase of the pulsar term. Fig. 2 shows the localization accuracy of GW sources as a function of the GW sky location under white noise levels of 10, 30, and 100 ns. In all cases, the accuracy is higher (the area is smaller) in regions where pulsars are densely distributed. It can be seen that even in the region with the best localization accuracy, $10^{-3}$ $\rm deg^{2}$ cannot be achieved. More details regarding the dependence of localization area on white noise level are provided in Appendix A.
@@
Figure 5: Left panel: The source numbers corresponding to the contours in Fig. 4. Right panel: Areas $\Omega$ of the two-dimensional 68% contours for GW sources located between J0437-4715 and J0030+0451. Source 14 is the same as Source 5. The green lines correspond to the fiducial case, while the blue lines correspond to Case 3. The solid, dashed, and dash-dotted lines correspond to the Slow, Fast, and No evolution cases in Table 1, respectively.
@@ VI Conclusions and Discussions
One of the limitations of this study is the assumption that the CGW signal can be described by a linearized model under the condition that the SNR is sufficiently high, referred to as the high-SNR/LSA limit. However, this assumption does not hold for current PTA observations, which remain in the low-SNR regime. Furthermore, recent analyses have shown that the posterior distribution can exhibit a multimodal structure in sky location [23], whereas the closed-form posterior distribution is unimodal and therefore cannot capture this feature.
[…]
Another limitation is that, although we performed Gaussian integration to obtain the closed-form expression, the parameters of stochastic signals cannot be treated as free parameters. Investigating the behavior of simultaneous parameter estimation of a CGW and stochastic signals remains a subject for future work. In particular, understanding the simultaneous parameter estimation of common-spectrum spatially uncorrelated red noise and a CGW, as demonstrated in [43], is likely to become increasingly important. Recent studies also emphasized the importance of resolving a CGW and the stochastic GW background simultaneously [24, 28].
@@ Acknowledgements.
We would like to thank Polina Petrov for helpful discussions. We are also grateful to the referee for valuable comments and suggestions. RK is supported by JSPS KAKENHI Grant Number 24K17051. KT is partially supported by JSPS KAKENHI Grant Numbers 20H00180, 21H01130, 21H04467, and 24H01813, Bilateral Joint Research Projects of JSPS, and the ISM Cooperative Research Program (2023-ISMCRP-2046).
Appendix A Dependence of area on white noise level
Fig. 6 shows how the estimated area depends on the white noise level. It can be seen that as the white noise level decreases, the result approaches that of the uniform prior case. Conversely, as the noise level increases, it approaches that of the fixed parameter case. The intermediate region between the uniform prior case and the fixed-parameter case is where the Gaussian prior exhibits nontrivial behavior. In the fiducial case, the difference between the Gaussian prior and the uniform prior becomes apparent when the white noise exceeds 10 ns.
$\propto\frac{1}{\sigma_{n}},$ (51)
where $\vec{\theta}_{\rm ML}$ is the maximum likelihood estimator. This relationship is consistent with the result of a frequentist study [52]. It is known that, under the high-SNR/LSA limit, the covariance matrix of the posterior distribution with a uniform prior, as shown in Eq. 39, coincides with that of the maximum likelihood estimator [49]. Since the area is characterized by the covariance matrix, this consistency is an expected result. It is also known that the mean of the maximum likelihood estimator equals the true value. Therefore, in Fig. 6, we approximated the SNR by substituting the true value in place of the estimator itself. This SNR, evaluated using the true parameters of the signal model, corresponds to the optimal SNR of the true signal as defined in Section IV.2. Although the difference between the maximum likelihood estimator and its mean is expected to increase with larger white noise, this effect is beyond the scope of the present analysis.
Finally, it should be noted that the high-SNR/LSA limit underlying the closed-form posterior distribution is valid only when the optimal SNR of the true signal is sufficiently high. If the SNR becomes too low, the resulting posterior deviates from the true distribution. For example, when the rms value of the white noise is comparable to the amplitude of the Earth term, the SNR is approximately 100 in Fig. 6, which is sufficiently high to justify the use of the high-SNR/LSA limit.
Appendix B Contour of the multivariate normal distribution
@@
Substituting Eq. 53 into Eq. 52 shows that the probability density on the surface of the ellipsoid is independent of $\vec{x}$ and is constant. This surface is called a contour. According to Eq. 53, it is represented as an ellipsoid centered at $\vec{\mu}$, with axes given by $\pm\sqrt{\chi^{2}_{k}(\alpha)\lambda_{i}}\vec{e}_{i}$, where ($\lambda_{i}$, $\vec{e}_{i}$) is an eigenvalue-eigenvector pair of $\Sigma$, which satisfies $\Sigma\vec{e}_{i}=\lambda_{i}\vec{e}_{i}$ for $i=1,2,\cdots,k$.